OQ Inc. is an all equity financed company in the country which does not have any corporate taxes. The company has an EBIT of $2 million and EBIT is expected to grow at 6% per year forever. The cost of equity for OQ Inc. is 18%. Currently, the company has 625,000 shares of common stock outstanding.

a. Calculate the value of the firm with its all equity financed capital structure.


b. Suppose the company is thinking about changing its capital structure. The company wants to have 35% debt and 65% equity financing. Suppose the company can borrow the required amount at 10% per year. The company is going to buy 35% of its shares back with the money it is borrowing. Determine the amount of money the company needs to borrow to achieve its desired capital structure.


c. Determine the EBIT that would make the company indifferent between 35% debt and 65% equity financing, and 100% equity financing.

d. Given the break-even EBIT you calculated in part (c) of this question and the EBIT the company has, indicate and briefly justify the capital structure that should be preferred by the firm.

e. Briefly discuss if the value of the firm with the capital structure you indicate in part (d) of this question would be higher, lower or the same as the value you calculated in part (a) of this question.

f. Calculate the cost of equity and the weighted average cost of capital for the firm with the capital structure you indicated in part (d) of this question.

g. Suppose today is two years later and the country decides to have a corporate tax rate of 40%. Suppose OQ Inc. stayed as an all equity financed firm for the past two years. Calculate the value of OQ Inc. when there is a 40% corporate tax.

h. Suppose OQ Inc. decides to issue $3 million worth of bonds and use this money to buy back some of its shares. The interest rate on these bonds will be 10% per year. Calculate the value of the OQ Inc. and value of its equity with this new capital structure when there are corporate taxes.

i. Briefly explain the reason for the difference in value of the firm you calculated in parts (g) and (h) of this question.

j. Calculate the cost of equity and the weighted average cost of capital the company will have with its new capital structure when there is 40% corporate tax rate
Answer #1

You have asked a question with multiple sub parts. I have addressed the first four sub parts of the question. Please post the balance sub parts in a lot of four, separately.

Part (a)

the value of the firm with its all equity financed capital structure = EBIT x (1 + g) / (Keu - g)

= 2 x (1 + 6%) / (18% - 6%) = $ 17.67 million

Part (b)

The amount of money the company needs to borrow to achieve its desired capital structure = Debt = 35% of the value = 35% x 17.67 = $ 6.18 million

Part (c)

EPS should remain same in both the cases:

100% equity financing case, EPS = EBIT / N

35% debt case, EPS = (EBIT - Interest) / N*

Interest = Debt x 10% = 6.18 x 10% = 0.618

N* = 65% x N = 65% x 625,000 = 406,250

Hence, EBIT / 625,000 = (EBIT - 0.618) / 406,250

Hence, 406,250 x EBIT = 625,000 x (EBIT - 0.618)

Hence, Break even EBIT = 625,000 x 0.618 / (625,000 - 406,250) = $ 1.77 million

Part (d)

The company current has an EBIT of $ 2 million > 1.77 million = the break even EBIT

Hence, the EPS of the firm will be higher in case of higher leverage. The firm will prefer the capital structure with 35% debt and 65% equity financing. EPS in this case = (EBIT - I) / N* = (2 - 0.618) x 1,000,000 / 406,250 = $ 3.40 per share

EPS in case of 100% equity finance = EBIT / N = 2 x 1,000,000 / 625,000 = $ 3.20 per share

The factor of 1,000,000 is to convert the $ million to simple $ in the numerator

The company has thus higher EPS in case of 35% debt finance, hence the firm will prefer the capital structure with 35% debt and 65% equity financing.